OpenStudy (anonymous):

Marci finished the 100-yard dash 1 4/5 seconds faster than Eva. If Eva finished in 14 7/10 seconds, how long did it take Marci to finish?

OpenStudy (anonymous):

@zzr0ck3r help me please!

OpenStudy (zzr0ck3r):

m= marcie e = eva m-(1 4/5) = e m-(9/5) = e let e=14 7/10 = 147/10 m-(9/5) = (147/10) solve for m

OpenStudy (anonymous):

@zzr0ck3r I divide?